A Farmer has 340 yards of fencing and wants to enclose a rectangular area of 7200 square yards. What dimensions should he use?
Answer by Alan3354(69443)

P = 2L+ 2W = 340
Area = L*W = 7200
----
L + W = 170
Find a pair of factors of 7200 with a sum of 170.